# The Growing Gains and Pains of Iterative Web Corpora Crawling: Insights from South Slavic CLASSLA-web 2.0 Corpora

### v1

@@ 5.2.  Small Overlap Between the Two Crawls
The comparison of CLASSLA-web 1.0 and 2.0 in terms of sizes has shown that the new edition of web corpora comprises more texts – up to two times as much as the 1.0 version. In this section, we examine the degree of overlap between the 1.0 and 2.0 versions of the CLASSLA-web corpora. This analysis aims to determine whether the growth in corpus size of the 2.0 version can be partially attributed to the inclusion of the majority of the 1.0 corpora within the 2.0 corpora. Additionally, this investigation provides us with insights into the degree of change in the content of South Slavic webs over a short period between the two observed crawls.
We measure the content overlap by identifying near-duplicated texts using MinHash over word 4-grams, with a similarity threshold set to 0.7, defined via manual validation. Surprisingly, despite the fact that only two years have passed between the two crawls, there is a small content overlap – only around 20% of texts from version 1.0 are present in CLASSLA-web 2.0 corpora.
In Table 2, we report the share of texts unique to each crawling iteration. On average, 82% of texts (comprising approximately 81% of words) in CLASSLA-web 2.0 corpora are new relative to the CLASSLA-web 1.0 corpora. These findings indicate a rapid turnover of web content: most content present in version 1.0 has disappeared in two years and the majority of the content in version 2.0 has appeared recently.
Considering also the changes in corpora sizes, shown in Figure 4, Bulgarian (bg) and Montenegrin (cnr) are the fastest-changing and expanding webs. The CLASSLA-web 2.0 corpora from these two national webs nearly doubled in size compared to the 1.0 versions, while retaining only 11–14% of content that was already present in the 1.0 corpora. In contrast, Croatian web changed the least, with 24% of texts in version 2.0 already being present in the version 1.0.
@@
The small content overlap also suggests that through iterative crawling over a two-year period, it is possible to collect large quantities of texts. Figure 5 shows the combined size of CLASSLA-web 1.0 and CLASSLA-web 2.0, with texts from version 1.0 excluded from version 2.0. The total size of the merged CLASSLA-web corpora is 57 million texts, with the Bulgarian corpus amounting to 20 million texts. These results demonstrate the effectiveness of iterative crawling for collecting large-scale text resources for less-resourced languages, and provide a strong motivation to continue with biannual crawling efforts.
@@ 5.2.1.  URL Overlap as a Quick Insight into Content Overlap
In the previous section, we assessed the extent of overlapping content between two corpora using the MinHash algorithm to detect near-duplicate texts. Since this method is computationally intensive, we explore faster alternatives for gaining initial insights into content overlap. One potential indicator of degree of content overlap is the overlap of URLs of web pages from which the texts were collected, as this information is included in the metadata for each text. Figure 6 shows a correlation between the percentage of shared URLs and the percentage of shared texts across both corpora. This relationship, based on seven pairs of national CLASSLA-web 1.0 and 2.0 corpora (e.g., CLASSLA-web.hr 1.0 and CLASSLA-web.hr 2.0), is strong, with a Pearson correlation coefficient of 0.908 and a p-value of 0.0047.
@@
As shown in Figure 6, this correlation can be modelled using a weighted linear regression, implemented in the scikit-learn Python library [^11] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LinearRegression.html. The model was trained on URL and text overlap data from the seven CLASSLA-web 2.0 corpora, specifically their overlap with the corresponding CLASSLA-web 1.0 corpora. The weights were based on the normalized sizes of the CLASSLA-web 2.0 corpora. When tested on CLASSLA-web 1.0 data (i.e., the percentage of overlapping texts and URLs with version 2.0), the model’s predictions deviate by 0.4–5 percentage points, and below 1.7 percentage points on larger corpora, namely Bulgarian, Croatian, Slovenian and Serbian. On version 2.0, the error ranges from 0.15 to 4 points, with deviations of 1 point or less for the larger corpora.
textOverlap=1.3476+1.2161URLOverlaptextOverlap=1.3476+1.2161*URLOverlap (1)
Equation 1 shows the function derived from the regression model, which can be used to estimate the likely percentage of shared texts between two corpora (textOverlap) based on the percentage of overlapping URLs (URLOverlap). This function offers a rough but useful approximation of content overlap between large web corpora, enabling quick comparisons when detecting near-duplicates directly would be too time-consuming or computationally expensive.
@@ 5.3.  Degradation of Web Content
While the comparison between versions 1.0 and 2.0 of the CLASSLA-web corpora revealed substantial gains in text content, manual inspection of the most frequent domains also showed a notable increase in low-quality or automatically generated content. As shown in Figure 7, a manual review of the Western Slavic CLASSLA-web 2.0 corpora found that many of the top 250 domains were problematic, accounting for 15% of the total number of texts before their removal.
@@ 6.  Conclusion
In this paper, we present the CLASSLA-web 2.0 corpus collection for South Slavic languages, a result of an ongoing biannual web crawling infrastructure. Although the corpora were collected from the same national domains just two years after the CLASSLA-web 1.0 collection (Ljubešić and Kuzman, 2024), the new datasets are significantly larger and contain mostly new content. On average, the 2.0 versions include around 50% more words and texts than the 1.0 versions. In total, the CLASSLA-web 2.0 corpora consist of 17 billion words and 38 million texts in seven South Slavic languages. The corpora are freely available in the JSONL and VERT formats from the CLARIN.SI repository (Kuzman Pungeršek et al., 2026) [^12] http://hdl.handle.net/11356/2079 and can be queried via the CLARIN.SI concordancers [^13] https://www.clarin.si/ske/. More information on the corpora development and accessibility is available on a website, dedicated to the CLASSLA-web collections (https://clarinsi.github.io/classla-web/).
@@
Since the CLASSLA-web 2.0 corpora were developed using the same pipeline as the previous version, CLASSLA-web 1.0, this enables longitudinal analyses of changes in South Slavic web content between the two crawls, that is, between 2021 and 2024. Most notably, there has been a significant rise in low-quality or automatically generated content. Manual inspection of the most frequent domains in CLASSLA-web 2.0 reveals that low-quality domains account for a much larger share of both texts and words compared to version 1.0 and appear far more often among the top-ranked domains. Specifically, an inspection of the top 250 domains identified a large number of bad domains which contributed 15% of the total number of texts prior to their removal from the cleaned corpora. These findings underline the growing importance of a manual validation of high-frequency domains to maintain web corpus quality, especially in light of the widespread availability of tools for automatic large-scale text generation.
To conclude, despite the growing pains related to maintaining high-quality web corpora that would be free from automatically-generated texts, web crawling remains a highly effective method for collecting large-scale text resources for less-resourced languages such as South Slavic languages. Due to the rapid turnover of online content, repeating the crawl every two years can more than double the size of the total collected corpora and our results show small content overlap between versions collected two years apart. Thanks to their size and recency, the CLASSLA-web 2.0 corpora have already attracted interest within the South Slavic research and language technology development communities. The corpora have been included in the pretraining data for Slovenian open-source large language models (Vreš et al., 2024) and introduced to linguists through recent editions of the CLASSLA-Express workshops that focus on the use of web corpora and concordancers for linguistic analyses (Ljubešić et al., 2024a). Highly motivated by the community’s uptake of this new web corpus collection, we will continue our biannual web crawling efforts to provide increasingly bigger and up-to-date web corpora for South Slavic languages.
@@ 8.  Acknowledgements
The research presented in this paper was conducted within the research project “Basic Research for the Development of Spoken Language Resources and Speech Technologies for the Slovenian Language” (J7-4642), the research project “Large Language Models for Digital Humanities” (GC-0002), the research project “Embeddings-based techniques for Media Monitoring Applications” (L2-50070, co-funded by the Kliping d.o.o. agency) and within the research programme “Language resources and technologies for Slovene” (P6-0411), all funded by the Slovenian Research and Innovation Agency (ARIS).

### v2

@@ 5.2.  Small Overlap Between the Two Crawls
A comparison of CLASSLA-web 1.0 and 2.0 shows that the new edition of the web corpora contains up to twice as many texts as version 1.0. In this section, we examine the degree of overlap between the 1.0 and 2.0 versions of the CLASSLA-web corpora. This analysis aims to determine whether the growth in corpus size of the 2.0 version can be partially attributed to the inclusion of the majority of the 1.0 corpora within the 2.0 corpora. Additionally, this investigation provides us with insights into the degree of change in the content of South Slavic webs over a short period between the two observed crawls.
We measure content overlap by identifying near-duplicated texts using MinHash over word 4-grams, with a similarity threshold set to 0.7, defined via manual validation. Surprisingly, despite the fact that only two years have passed between the two crawls, there is a small content overlap – only around 20% of texts from version 1.0 are present in the CLASSLA-web 2.0 corpora. [^11] Detailed information indicating which texts in version 2.0 are near duplicates of specific texts in version 1.0 is provided at https://github.com/clarinsi/classla-web/tree/main/more-info/duplicates-with-1.0.
In Table 2, we report the share of texts unique to each crawling iteration. On average, 82% of texts (comprising approximately 81% of words) in CLASSLA-web 2.0 corpora are new relative to the CLASSLA-web 1.0 corpora. These findings indicate a rapid turnover of web content: most of the content present in version 1.0 has disappeared in two years, and the majority of the content in version 2.0 has appeared recently.
Taking into account also changes in corpora sizes, shown in Figure 4, Bulgarian (bg) and Montenegrin (cnr) are the fastest-changing and expanding webs. The CLASSLA-web 2.0 corpora from these two national webs nearly doubled in size compared to the 1.0 versions, while retaining only 11–14% of content that was already present in the 1.0 corpora. In contrast, Croatian web changed the least, with 24% of texts in version 2.0 already being present in version 1.0.
@@
The small content overlap also suggests that through iterative crawling over a two-year period, it is possible to collect large quantities of texts. Figure 5 shows the combined size of CLASSLA-web 1.0 and CLASSLA-web 2.0, with texts from version 1.0 excluded from version 2.0. The total size of the merged CLASSLA-web corpora is 57 million texts, with the Bulgarian corpus amounting to 20 million texts. These results demonstrate the effectiveness of iterative crawling for collecting large-scale text resources for less-resourced languages, and provide a strong motivation to continue with biennial crawling efforts.
@@ 5.2.1.  URL Overlap as a Quick Insight into Content Overlap
In the previous section, we assessed the extent of overlapping content between two corpora using the MinHash algorithm to detect near-duplicate texts. Since this method is computationally intensive, we explore faster alternatives for gaining initial insights into content overlap. One potential indicator of the degree of content overlap is the overlap of URLs of web pages from which the texts were collected, as this information is included in the metadata for each text. Figure 6 shows a correlation between the percentage of shared URLs and the percentage of shared texts across both corpora. This relationship, based on seven pairs of national CLASSLA-web 1.0 and 2.0 corpora (e.g., CLASSLA-web.hr 1.0 and CLASSLA-web.hr 2.0), is strong, with a Pearson correlation coefficient of 0.908 and a p-value of 0.0047.
@@
As shown in Figure 6, this correlation can be modelled using a weighted linear regression, implemented in the scikit-learn Python library [^12] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LinearRegression.html. The model was trained on URL and text overlap data from the seven CLASSLA-web 2.0 corpora, specifically their overlap with the corresponding CLASSLA-web 1.0 corpora. The weights were based on the normalised sizes of the CLASSLA-web 2.0 corpora. When tested on CLASSLA-web 1.0 data (i.e., the percentage of overlapping texts and URLs with version 2.0), the model’s prediction errors are 0.4–5 percentage points, and below 1.7 percentage points on larger corpora, namely Bulgarian, Croatian, Slovenian and Serbian. On version 2.0, the error ranges from 0.15 to 4 points, with deviations of 1 point or less for the larger corpora.
$textOverlap=1.347+1.216\cdot URLOverlap$ (1)
Equation 1 shows the function derived from the regression model, which can be used to estimate the likely percentage (0-100) of shared texts between two corpora (textOverlap) based on the percentage of overlapping URLs (URLOverlap). This function offers a rough but useful approximation of content overlap between large web corpora, enabling quick comparisons when detecting near-duplicates directly would be too time-consuming or computationally expensive.
@@ 5.3.  Degradation of Web Content
While the comparison between versions 1.0 and 2.0 of the CLASSLA-web corpora revealed substantial gains in text content, manual inspection of the most frequent domains also showed a notable increase in low-quality or automatically generated content. [^13] Lists of domains from the CLASSLA-web corpora that have been manually verified or identified as low quality are provided at https://github.com/clarinsi/classla-web/tree/main/urls. As shown in Figure 7, a manual review of the Western South Slavic CLASSLA-web 2.0 corpora found that many of the top 250 domains were problematic, accounting for 15% of the total number of texts before their removal.
@@ 6.  Conclusion
In this paper, we present the CLASSLA-web 2.0 corpus collection for South Slavic languages, a result of an ongoing biennial web crawling infrastructure. Although the corpora were collected from the same national domains just two years after the CLASSLA-web 1.0 collection (Ljubešić and Kuzman, 2024), the new datasets are significantly larger and contain mostly new content. On average, the 2.0 versions include around 50% more words and texts than the 1.0 versions. In total, the CLASSLA-web 2.0 corpora consist of 17 billion words and 38 million texts in seven South Slavic languages. The corpora are freely available in JSONL and VERT formats from the CLARIN.SI repository (Kuzman Pungeršek et al., 2026) [^14] http://hdl.handle.net/11356/2079 and can be queried via the CLARIN.SI concordancers [^15] https://www.clarin.si/ske/. More information on corpora development and accessibility is available on a website dedicated to the CLASSLA-web collections (https://clarinsi.github.io/classla-web/).
@@
Since the CLASSLA-web 2.0 corpora were developed using the same pipeline as the previous version, CLASSLA-web 1.0, this enables longitudinal analyses of changes in South Slavic web content between the two crawls, that is, between 2021 and 2024. Most notably, there has been a significant increase in low-quality or automatically generated content. Manual inspection of the most frequent domains in CLASSLA-web 2.0 reveals that low-quality domains account for a much larger share of both texts and words compared to version 1.0 and appear far more often among the top-ranked domains. Specifically, an inspection of the top 250 domains identified a large number of bad domains, which contributed 15% of the total number of texts prior to their removal from the cleaned corpora. These findings underline the growing importance of a manual validation of high-frequency domains to maintain web corpus quality, especially in light of the widespread availability of tools for automatic large-scale text generation.
To conclude, despite the growing pains related to maintaining high-quality web corpora that would be free from automatically-generated texts, web crawling remains a highly effective method for collecting large-scale text resources for less-resourced languages such as South Slavic languages. Due to the rapid turnover of online content, repeating the crawl every two years can more than double the size of the total collected corpora and our results show small content overlap between versions collected two years apart. Thanks to their size and recency, the CLASSLA-web 2.0 corpora have already attracted interest within the South Slavic research and language technology development communities. The corpora have been included in the pretraining data for Slovenian open-source large language models (Vreš et al., 2024) and introduced to linguists through recent editions of the CLASSLA-Express workshops that focus on the use of web corpora and concordancers for linguistic analyses (Ljubešić et al., 2024a). Highly motivated by the community’s uptake of this new web corpus collection, we will continue our biennial web crawling efforts to provide increasingly bigger and up-to-date web corpora for South Slavic languages.
@@ 8.  Acknowledgements
The research presented in this paper was conducted within the research project “Basic Research for the Development of Spoken Language Resources and Speech Technologies for the Slovenian Language” (J7-4642), the research project “Large Language Models for Digital Humanities” (GC-0002), the research project “Embeddings-based techniques for Media Monitoring Applications” (L2-50070, co-funded by the Kliping d.o.o. agency), the Research Infrastructure DARIAH-SI (I0-E007), and within the research programme “Language resources and technologies for Slovene” (P6-0411), all funded by the Slovenian Research and Innovation Agency (ARIS).
This research was also supported by LLMs4EU, co-funded by the Digital Europe Programme under GA 101198470. This project is co-funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@ 9.  Bibliographical References
10.  Language Resource References
[…]
  • Kuzman and Ljubešić (2023) Kuzman, Taja and Ljubešić, Nikola. 2023. Multilingual text genre classification model X-GENRE. Hugging Face. PID 10.57967/hf/0927.
  • Kuzman and Ljubešić (2024) Kuzman, Taja and Ljubešić, Nikola. 2024. Multilingual text genre classification model X-GENRE. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1961.
  • Kuzman and Ljubešić (2025) Kuzman, Taja and Ljubešić, Nikola. 2025. Multilingual IPTC News Topic Classifier. Hugging Face. PID 10.57967/hf/4709.
  • Kuzman Pungeršek et al. (2026) Kuzman Pungeršek, Taja and Rupnik, Peter and Ljubešić, Nikola. 2026. South Slavic web corpus collection CLASSLA-web 2.0. Slovenian Language Resource Repository CLARIN.SI. PID http://hdl.handle.net/11356/2079.
  • Ljubešić et al. (2024a) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024a. Bosnian web corpus CLASSLA-web.bs 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1927.
  • Ljubešić et al. (2024b) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024b. Bulgarian web corpus CLASSLA-web.bg 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1928.
  • Ljubešić et al. (2024c) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024c. Croatian web corpus CLASSLA-web.hr 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1929.
  • Ljubešić et al. (2024d) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024d. Macedonian web corpus CLASSLA-web.mk 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1932.
  • Ljubešić et al. (2024e) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024e. Montenegrin web corpus CLASSLA-web.cnr 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1930.
  • Ljubešić et al. (2024f) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024f. Serbian web corpus CLASSLA-web.sr 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1931.
  • Ljubešić et al. (2024g) Ljubešić, Nikola and Rupnik, Peter and Kuzman, Taja. 2024g. Slovenian web corpus CLASSLA-web.sl 1.0. Slovenian language resource repository CLARIN.SI. PID http://hdl.handle.net/11356/1882.
  • Sites (2013) Sites, Dick. 2013. Compact language detector 2. GitHub. Software available at https://github.com/CLD2Owners/cld2 (last updated on August 2015).